# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v5

@@ Abstract
This work proposes a unified framework for portfolio allocation, covering both asset selection and optimization, based on a multiple-hypothesis predict-then-optimize approach. The portfolio is modeled as a structured ensemble, where each predictor corresponds to a specific asset or hypothesis. Structured ensembles formally link predictors’ diversity — captured via ensemble loss decomposition — to out-of-sample risk diversification. A structured dataset of predictor outputs is constructed with a parametric diversity control, influencing both the training process and diversification outcomes. This dataset feeds a supervised ensemble model whose target portfolio must align with the ensemble combiner rule implied by the loss. For squared loss, the arithmetic mean applies, yielding the equal-weighted portfolio as the optimal target. For asset selection, a novel method is introduced that prioritizes assets from more diverse predictor sets, even at the cost of lower average predicted returns, via a diversity–quality trade-off. This diversity is applied prior to optimization and can integrate into other allocation techniques. Experiments on SP500 stocks validate the theoretical framework and show that both sources of diversity expand the limits of portfolio diversification, achieving strong performance across one-step and multi-step allocation tasks.
@@ 1 Introduction
Few researchers have also employed the use of predictive models to optimize decision-making processes, often with the objective of minimizing decision-related costs. Such conventional ”predict-then-optimize” paradigm decouples prediction and decision-making into two distinct stages: first, a predictive model is developed to maximize predictive accuracy; second, decisions are derived based on the model’s outputs and associated cost functions. A significant limitation of this approach is its failure to incorporate cost considerations during the predictive modeling phase (Donti and Kolter, 2017). To address this shortcoming, recent advances have introduced integrated methodologies such as ”predict-and-optimize” (Wilder et al., 2019; Vanderschueren et al., 2022) and Smart ”Predict, then Optimize” (SPO) (Elmachtoub and Grigas, 2017). These approaches bridges the gap between structured prediction and cost-sensitive optimization, offering a robust solution for enhancing decision-making processes (Kotary et al., 2023). These approaches, however, are computationally expensive and mostly focus on solving the classification problems (Goh and Jaillet, 2016).
@@
While the aforementioned solutions reduce parameter uncertainty and can improve out-of-sample performance, they fail to establish a connection between in-sample and out-of-sample diversification, which is a key feature addressed in this work. In this context, a framework is proposed for portfolio optimization using structured ensembles for prediction with multiple hypotheses. Each portfolio component corresponds to a hypothesis, with its return series modeled by an individual predictor. This approach differs from return forecasting (Ma et al., 2021) or scoring models (Nguyen and Lo, 2012; Zhao et al., 2022), where models are used to score assets and predict their returns before applying optimization techniques for portfolio allocation. Building on prior research, the degree of diversity in ensemble learning is parametrically controlled during predictor training and encoded in a structured data set of predictions (Dominguez et al., 2025). This dataset serves as input for an ensemble model designed to predict portfolio returns. The ensemble model is optimized as a supervised prediction problem, using as a target a portfolio whose weights are chosen to match the ensemble combiner rule derived from the Bias-Variance-Diversity decomposition (Wood et al., 2024). In the case of squared loss, the corresponding combiner is the arithmetic mean, implying that the equal-weighted portfolio should be selected. The optimal portfolio weights are then obtained by normalizing the optimal ensemble parameters. To the best of the authors’ knowledge, this is the first plug-in method that enables parametrically controlled portfolio risk diversification prior to the optimization step.
@@
Additionally, this work also investigates the hypothesis that out-of-sample portfolio diversification relates to the asset or hypothesis selection. Diversity in asset selection is controlled parametrically, independently of the portfolio optimization phase. A parameter regulates the diversity of asset sets—randomly selected from ranked out-of-sample return predictions—across various portfolio sizes and parameters. Results show that greater diversity in asset sets increases out-of-sample portfolio diversification and risk-adjusted returns, even when the sets have lower average returns. This work links both sources of diversity (i.e., asset selection and portfolio weight optimization) to out-of-sample portfolio diversification and demonstrates that the diversification levels from both sources can be parametrically controlled prior to decision-making, offering valuable insights for practitioners.
@@
    Introduced a consistent framework for applying structured ensemble models in multi-hypothesis prediction settings for portfolio optimization, where the ensemble combiner rule defines the portfolio target. Specifically, in the case of squared loss, this corresponds to the equal-weighted portfolio.
@@
    Presented a ”plug-in” and robust method enabling parametric a priori control over out-of-sample portfolio risk diversification by linking ensemble diversity, derived from the Bias-Variance-Diversity loss decomposition (Wood et al., 2024), to portfolio risk diversification.
@@
    Proposed a strategy for parametrically controlling diversity in asset or hypothesis selection, leveraging a diversity–quality trade-off in predicted returns to enhance out-of-sample portfolio diversification.
@@
    Systematically validated the expansion of diversification limits (Dalio, 2018) enabled by the proposed framework, as well as through the incorporation of diversity in the asset selection stage across other methodologies, demonstrated empirically in both one-step and multi-step decision settings.
The paper is organized as follows. Section 2 provides a review of previous work on data-driven and robust optimization, Predict-then-Optimize frameworks, structured ensemble learning, and diversity. Section 3 describes the proposed framework and methodology. Section 4 presents the experimental results and discussions. Finally, Section 5 offers the final remarks and future outlook.
@@ 2 Literature Review
Portfolio optimization solutions can be broadly categorized into data-driven approaches and robust optimization methods. Data-driven methods typically rely on plug-in optimization, where the objective function is predefined by frameworks such as Modern Portfolio Theory (MPT), and data is used solely to estimate model parameters. In contrast, robust optimization modifies the objective function and relaxes certain framework assumptions to improve adaptability under uncertainty. This approach was among the first to incorporate loss functions, which were later extended to predictive models and predict-then-optimize frameworks, albeit with computational challenges. Ensemble-based methods further introduce loss functions that account for diversity in predictions. This work integrates ensemble models with multiple hypothesis prediction, establishing a connection between diversity in learning and risk diversification, while also introducing computationally efficient analytical solutions.
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches leverage historical and real-time market data to model uncertainties, employing methods such as stochastic programming, machine learning, and empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022; Ma et al., 2021; Ozelim et al., 2023). Robust portfolio optimization solutions include fuzzy approaches (Wu and Liu, 2012), genetic algorithms (Akopov, 2014), multiobjective particle swarm optimization (Chen and sha Zhou, 2018), and learning-guided multiobjective evolutionary algorithms (Lwin et al., 2014). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thus addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 2.2 Predictive Modeling Integrated with Decision Optimization
2.3 Diversity in Ensemble Learning and Portfolio Diversification
@@
Researchers link portfolio estimation to statistical decision theory by treating the difference in investor utility between the ”true” optimal portfolio and the plug-in portfolio as an economic loss function (Kan and Zhou, 2007). Outcomes are enhanced by incorporating the investor’s utility objective directly into the statistical weight estimation process like in Bayesian decision theoretic approaches, rather than addressing estimation and utility maximization as separate issues (Avramov and Zhou, 2010; Kan and Zhou, 2007). Research has demonstrated the economic implications of using Mean Square Error (MSE) in portfolio optimization and financial forecasting, showing that integrating MSE into decision-making processes can enhance portfolio performance by minimizing estimation errors and optimizing risk-return trade-offs (Lasse Heje Pedersen and Levine, 2021; Mahadi et al., 2022; Cai et al., 2024). This work extends this idea by incorporating diversity into the MSE loss function within an ensemble learning framework. Furthermore, it establishes for the first time a connection between this notion of diversity and portfolio diversification, proposing parametric methods to manage diversification prior to decision-making.
@@ 3.1 Preliminary
If the aforementioned portfolio optimization can be formulated as a prediction task, then the objective utility function of the mean-variance in (1) can be represented as an MSE loss function (Cai et al., 2024). This setup allows for finding the optimal weights that minimize the total prediction error of the portfolio, accounting for both bias and variance, while ensuring that the expected return remains above the target value. In order to incorporate the diversity into such an optimization, this work applies ensemble models to harness diversity in ensemble learning and linking it to portfolio diversification. In a multiple-hypothesis prediction setup in which each hypothesis is focused on a particular portfolio constituent, the portfolio can be model as an ensemble of hypotheses so that the diversity term in the Bias-Variance-Diversity decomposition with the MSE Loss can be connected to diversification and controlled parametrically, while the centroid combiner rule is respected in the portfolio target selection. To elaborate, in a supervised setting with the data set of time series returns ($\boldsymbol{x}$, $\boldsymbol{r}$), the Bias-variance-diversity decomposition can be expressed in terms of the MSE loss $\mathcal{L}(\boldsymbol{r},\boldsymbol{\bar{r}})$ as (Wood et al., 2024):
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
where $g(\cdot)$ denotes the centroid combiner rule (e.g., averaging, voting, or weighted combination) for the ensemble (Wood et al., 2024). Individual predictors are trained with stochastic gradient descent (SGD) with a parameter $\varepsilon\in[0,1]$ that controls their diversity in learning at each step $i$ so that, for the $j$-th predictor, the parameter update rule follows:
@@
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the output of the $j$-th predictor is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the predictors that perform the best, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble. Algorithm 1 provides the pseudo-code for the structured dataset formation with stochastic gradient descent.
@@ 3.3 Predict-then-optimize Method with PSEM
Figure 1 shows a representation of the PSEM optimization, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $\varepsilon$. Once the training of the individual predictors is completed, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the Structured Radial Basis Function Network (s-RBFN) (Dominguez et al., 2025), shown in this figure and trained using least squares.
@@ 3.4 Example with Structured Radial Basis Function Network (s-RBFN)
In this case, the structured dataset is used as input to a radial basis function network, where each $j$-th predictor $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$ is associated with a specific basis function $\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. The basis function $\phi_{j}$ is applied elementwise to the entries of the $j$-th input vector of predictions $\boldsymbol{\hat{x}}_{j}(\varepsilon)\in\mathbb{R}^{N}$, with $\mu_{j}$ and $\sigma_{j}$ representing the center and scale parameters specific to the $j$-th predictor. This defines a transformation map $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon)):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$, where $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon))_{ij}=\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. For example, the Gaussian basis function $\phi_{j}\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{\hat{x}}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ can be used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{\hat{x}}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
@@
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from Modern Portfolio Theory (MPT) (Markowitz, 1952), shown in the blue box on the left, and the s-RBFN model trained via least squares (Dominguez et al., 2025), as the PSEM model, shown in the blue box on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Both approaches allow closed-form solutions via least squares, rendering them computationally equivalent in terms of efficiency. In the PSEM framework, the input data $\boldsymbol{x}$ is replaced with the diversified structured dataset of predictions $\boldsymbol{\hat{x}}(\varepsilon)$, which is then augmented using basis functions within the s-RBFN model. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
The focus of this work is on portfolio allocation, which encompasses both asset selection and portfolio optimization. The latter has been addressed in the previous sections. The former is introduced in the next section through a novel method that leverages the diversity of asset return predictions as a asset selection tool. This approach seeks to uncover connections between diversity in asset selection and out-of-sample portfolio diversification, identify new limits on portfolio diversification introduced by this additional source of variability, and evaluate the ability to control these effects parametrically prior to weight optimization.

### v6

@@ Abstract
This work proposes a unified framework for portfolio allocation, covering both asset selection and optimization, based on a multiple-hypothesis predict-then-optimize approach. The portfolio is modeled as a structured ensemble, where each predictor corresponds to a specific asset or hypothesis. Structured ensembles formally link predictors’ diversity, captured via ensemble loss decomposition, to out-of-sample risk diversification. A structured data set of predictor output is constructed with a parametric diversity control, which influences both the training process and the diversification outcomes. This data set is used as input for a supervised ensemble model, the target portfolio of which must align with the ensemble combiner rule implied by the loss. For squared loss, the arithmetic mean applies, yielding the equal-weighted portfolio as the optimal target. For asset selection, a novel method is introduced which prioritizes assets from more diverse predictor sets, even at the expense of lower average predicted returns, through a diversity–quality trade-off. This form of diversity is applied before the portfolio optimization stage and is compatible with a wide range of allocation techniques. Experiments conducted on the full S&P 500 universe and a data set of 1,300 global bonds of various types over more than two decades validate the theoretical framework. Results show that both sources of diversity effectively extend the boundaries of achievable portfolio diversification, delivering strong performance across both one-step and multi-step allocation tasks.
@@ 1 Introduction
Some researchers have also employed the use of predictive models to optimize decision-making processes, often with the objective of minimizing decision-related costs. Such a conventional ”predict-then-optimize” paradigm decouples prediction and decision-making into two distinct stages: firstly, a predictive model is developed to maximize predictive accuracy; secondly, decisions are derived based on the model’s outputs and associated cost functions. A significant limitation of this approach is its failure to incorporate cost considerations during the predictive modeling phase (Donti and Kolter, 2017). To address this shortcoming, recent advances have introduced integrated methodologies such as ”predict-and-optimize” (Wilder et al., 2019; Vanderschueren et al., 2022) and Smart ”Predict, then Optimize” (SPO) (Elmachtoub and Grigas, 2017). These approaches bridge the gap between structured prediction and cost-sensitive optimization, offering a robust solution for enhancing decision-making processes (Kotary et al., 2023). These approaches, however, are computationally expensive and mostly focus on solving the classification problems (Goh and Jaillet, 2016).
@@
While the aforementioned solutions reduce parameter uncertainty and can improve out-of-sample performance, they fail to establish a connection between in-sample and out-of-sample diversification, which is a key feature addressed in this work. In this context, a framework is proposed for portfolio optimization using structured ensembles for prediction with multiple hypotheses. Each portfolio component corresponds to a hypothesis, with its return series modeled by an individual predictor. This approach differs from return forecasting (Ma et al., 2021) or scoring models (Nguyen and Lo, 2012; Zhao et al., 2022), where models are used to score assets and predict their returns before applying optimization techniques for portfolio allocation. Building on prior research, the degree of diversity in ensemble learning is parametrically controlled during predictor training and encoded in a structured data set of predictions (Dominguez et al., 2025). This data set serves as input for an ensemble model designed to predict portfolio returns. The ensemble model is optimized as a supervised prediction problem, using as a target a portfolio whose weights are chosen to match the ensemble combiner rule derived from the Bias-Variance-Diversity decomposition (Wood et al., 2024). In the case of squared loss, the corresponding combiner is the arithmetic mean, implying that the equal-weighted portfolio should be selected. The optimal portfolio weights are then obtained by normalizing the optimal ensemble parameters. To the best of the authors’ knowledge, this is the first method which connects ensemble learning diversity with portfolio risk diversification, enabling its parametric control before the optimization stage.
@@
Additionally, this work also investigates the hypothesis that out-of-sample portfolio diversification relates to the asset or hypothesis selection. Diversity in asset selection is controlled parametrically, independently of the learning phase. A parameter regulates the diversity of asset sets—randomly selected from ranked out-of-sample return predictions—across various portfolio sizes and parameters. Results show that greater diversity in asset sets increases out-of-sample portfolio diversification and risk-adjusted returns, even when the sets have lower average returns. This work links both sources of diversity (i.e., asset selection and portfolio weight optimization) to out-of-sample portfolio diversification and demonstrates that the diversification levels from both sources can be parametrically controlled prior to decision-making, offering valuable insights for practitioners.
@@
    A consistent framework is introduced for applying structured ensemble models in multi-hypothesis prediction settings for portfolio optimization, where the ensemble combiner rule defines the portfolio target. Specifically, in the case of squared loss, this corresponds to the equal-weighted portfolio.
@@
    It presents a ”plug-in” and robust method enabling parametric a priori control over out-of-sample portfolio risk diversification by linking ensemble diversity, derived from the Bias-Variance-Diversity loss decomposition (Wood et al., 2024), to portfolio risk diversification.
@@
    A strategy is proposed for parametrically controlling diversity in asset or hypothesis selection, leveraging a diversity–quality trade-off in predicted returns to enhance out-of-sample portfolio diversification.
@@
    The expansion of diversification limits (Dalio, 2018) is validated by the proposed framework, as well as through the incorporation of diversity in the asset selection stage across other methodologies, demonstrated empirically in both one-step and multi-step decision settings.
The paper is organized as follows: Section 2 provides a review of previous work on data-driven and robust optimization, Predict-then-Optimize frameworks, structured ensemble learning, and diversity. Section 3 describes the proposed framework and methodology. Section 4 presents the experimental results and discussions. Finally, in Section 5 the final remarks and future outlook are offered.
@@ 2 Literature Review
Portfolio optimization solutions can be broadly categorized into data-driven approaches and robust optimization methods. Data-driven methods typically rely on plug-in optimization, where the objective function is predefined by frameworks such as Modern Portfolio Theory (MPT), and data is used solely to estimate model parameters. In contrast, robust optimization modifies the objective function and relaxes certain framework assumptions to improve adaptability while tackling uncertainty. This approach was among the first to incorporate loss functions, which were later extended to predictive models and predict-then-optimize frameworks, albeit with computational challenges. Ensemble-based methods further introduce loss functions which account for diversity in predictions. This work integrates ensemble models with multiple hypothesis prediction, establishing a connection between diversity in learning and risk diversification, while also introducing computationally efficient analytical solutions.
@@ 2.1 Data-Driven and Robust Optimization
Reducing uncertainty in the optimization stage is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches leverage historical and real-time market data to model uncertainties, employing methods such as stochastic programming, machine learning, and empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022; Ma et al., 2021; Ozelim et al., 2023). Robust portfolio optimization solutions include fuzzy approaches (Wu and Liu, 2012), genetic algorithms (Akopov, 2014), multiobjective particle swarm optimization (Chen and sha Zhou, 2018), and learning-guided multiobjective evolutionary algorithms (Lwin et al., 2014). Building on these, this work presents a mixed method which combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization. Therefore, this addresses the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 2.2 Predictive Modeling Integrated with Decision Optimization
2.3 Diversity in Multi-Hypothesis Learning Ensembles for Portfolio Diversification
@@
Researchers link portfolio estimation to statistical decision theory by treating the difference in investor utility between the ”true” optimal portfolio and the plug-in portfolio as an economic loss function (Kan and Zhou, 2007). Outcomes are enhanced by incorporating the investor’s utility objective directly into the statistical weight estimation process as in Bayesian decision theoretic approaches, rather than addressing estimation and utility maximization as separate issues (Avramov and Zhou, 2010; Kan and Zhou, 2007). Research has demonstrated the economic implications of using Mean Square Error (MSE) in portfolio optimization and financial forecasting, showing that integrating MSE into decision-making processes can enhance portfolio performance by minimizing estimation errors and optimizing risk-return trade-offs (Lasse Heje Pedersen and Levine, 2021; Mahadi et al., 2022; Cai et al., 2024). This work extends this idea by incorporating diversity into the MSE loss function within an ensemble learning framework. Furthermore, for the first time, it establishes a connection between this notion of diversity and portfolio diversification, proposing parametric methods to manage diversification prior to decision-making.
@@ 3.1 Preliminary
If the aforementioned portfolio optimization can be formulated as a prediction task, then the objective utility function of the mean-variance in (1) can be represented as an MSE loss function (Cai et al., 2024). This setup allows for finding the optimal weights which minimize the total prediction error of the portfolio, accounting for both bias and variance, while ensuring that the expected return remains above the target value. In order to incorporate diversity into such an optimization, this work applies ensemble models to harness diversity in ensemble learning and link it to portfolio diversification.
In a multiple-hypothesis prediction setup in which each hypothesis is focused on a particular portfolio constituent, the portfolio can be modeled as an ensemble of hypotheses. Following that, the diversity term in the Bias-Variance-Diversity decomposition with the MSE Loss can be connected to diversification and controlled parametrically, while the centroid combiner rule is applied as the portfolio target. To elaborate, in a supervised setting with the data set of time series returns ($\boldsymbol{x}$, $\boldsymbol{r}$), the Bias-variance-diversity decomposition can be expressed in terms of the MSE loss $\mathcal{L}(\boldsymbol{r},\boldsymbol{\bar{r}})$ as (Wood et al., 2024):
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
where $g(\cdot)$ denotes the centroid combiner rule (e.g., averaging, voting, or weighted combination) for the ensemble (Wood et al., 2024). A diversity parameter $\varepsilon\in[0,1]$ modulates each predictor’s update at step $i$, such that the update rule for the $j$-th predictor is given by:
@@
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the output of the $j$-th predictor is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the predictors which perform the best, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble. Algorithm 1 provides the pseudo-code for the structured data set formation with stochastic gradient descent.
@@ 3.3 Predict-then-optimize Method with PSEM
Although the ensemble prediction $\bar{\boldsymbol{r}}=\frac{1}{M}\sum_{j=1}^{M}\hat{\boldsymbol{x}}_{j}(\varepsilon)$ is computed as an equally weighted average of base learner outputs, the optimal ensemble weights $\boldsymbol{w}$ derived from (8) do not generally coincide with uniform or equal-weighting. This discrepancy arises due to induced diversity via the parameter $\varepsilon$, variation in base learner initialization and regularization, and inherent dependencies among predictors. These factors collectively break the symmetry of the hypothesis space, resulting in non-uniform projections of the ensemble mean onto the span of predictions. As a result, the final ensemble weights adapt to capture these structural and informational asymmetries, ultimately improving loss alignment while diverging from naive equal weighting.
Figure 1 shows a representation of the PSEM optimization, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $\varepsilon$. Once the training of the individual predictors is completed, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the Structured Radial Basis Function Network (s-RBFN) (Dominguez et al., 2025), shown in this figure and trained using least-squares.
@@ 3.4 Example with Structured Radial Basis Function Network (s-RBFN)
In this case, the structured data set is used as input to a radial basis function network, where each $j$-th predictor $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$ is associated with a specific basis function $\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. The basis function $\phi_{j}$ is applied element-wise to the entries of the $j$-th input vector of predictions $\boldsymbol{\hat{x}}_{j}(\varepsilon)\in\mathbb{R}^{N}$, with $\mu_{j}$ and $\sigma_{j}$ representing the center and scale parameters specific to the $j$-th predictor. This defines a transformation map $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon)):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$, where $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon))_{ij}=\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. For example, the Gaussian basis function $\phi_{j}\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{\hat{x}}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ can be used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{\hat{x}}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
@@
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from Modern Portfolio Theory (MPT) (Markowitz, 1952), shown in the blue box on the left, and the s-RBFN model trained via the least-squares algorithm (Dominguez et al., 2025), as the PSEM model, shown in the blue box on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Both approaches allow closed-form solutions via the least-squares algorithm, rendering them computationally equivalent in terms of efficiency. In the PSEM framework, the input data $\boldsymbol{x}$ is replaced with the diversified structured data set of predictions $\boldsymbol{\hat{x}}(\varepsilon)$, which is then augmented using basis functions within the s-RBFN model. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
The focus of this work is on portfolio allocation, which encompasses both asset selection and portfolio optimization. The latter has been addressed in the previous sections. The former is introduced in the next section through a novel method which leverages the diversity of asset return predictions as an asset selection tool. This approach seeks to uncover connections between diversity in asset selection and out-of-sample portfolio diversification, identify new limits on portfolio diversification introduced by this additional source of variability, and evaluate the ability to control these effects parametrically prior to weight optimization.